OpenStudy (anonymous):

How do you: Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = x2 + 13

OpenStudy (anonymous):

x^2

OpenStudy (mathmate):

The general transformation of a quadratic is \( y=a(x-h)^2+k \) where h=horizontal translation to the right k= vertical translation upwards a=scaling factor. Can you proceed from here?

OpenStudy (anonymous):

So it would be k=13. So it would shift UP 13 units? down right left....?

OpenStudy (mathmate):

Yes, from y=x^2 to y=x^2+13, the graph is shifted upwards by 13 units by adding 13 to y.

OpenStudy (mathmate):

Down would be k=-13 Right would have been h=13, and left would have been h=-13.